Grade 1210 grade maths

Find the LCM and HCF of the following integers by applying the prime factorisation method.
12, 15 and 21

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

Writing the product of prime factors for all the three numbers,
we get,

12=2×2×3
15=5×3
21=7×3

Therefore,
HCF(12,15,21) = 3
LCM(12,15,21) = 2× 2 × 3 × 5 ×7 = 420

Thanks